OpenStudy (anonymous):

verify that is an indentity Sin(x-y)=(1-cotx tany)/(csc x sec y)

myininaya (myininaya):

\[\frac{1-\cot(x)\tan(y)}{\csc(x) \sec(y)}=\frac{1-\frac{\cos(x)}{\sin(x)} \frac{\sin(y)}{\cos(y)}}{\frac{1}{\sin(x)} \frac{1}{\cos(y)}} \cdot \frac{\sin(x)\cos(y)}{\sin(x)\cos(y)}\] \[=\frac{\sin(x)\cos(y)-\cos(x)\sin(y)}{1}\] you got it from here right?

OpenStudy (anonymous):

nope

myininaya (myininaya):

good thing i wasn't asking you then lol

OpenStudy (anonymous):

what do i do with the "1" in the denominator?

myininaya (myininaya):

a/1=a

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

good job,ty